Show that the dimensions of the displacement current ε^0dφ/dt are that of an electric current.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. ϵ base 0 d∅ base E/dt = ϵ base 0 EA/dt 4π ε base 0r^2 = M^-1L^-3T^4 A^2/M6-1L^-3 A^2 * A^1T^1/L^2 * L^2/T = A^1 = (Current) (proved).